OpenStudy (anonymous):

Write an equation of the line that passes through the point (1,2) and has the slope, m = 2.

OpenStudy (shamil98):

y = mx + b input your values and solve for b. 1 = 2(1) + b b = ? after finding b, replace it in this : y = 2x + b